a garden has a fence that is 48 feet long. if the width of the rectangular garden is 13 feet, what is the length of the garden?
Answer by ReadingBoosters(3246)

fence is perimeter = 48 = 2(l+w) = 2l + 2w
w = 13
2l + 2(13) = 48
2l + 26 = 48
2l + 26 - 26 = 48 - 26
2l = 22
2l/2=22/2
l=11